OpenStudy (anonymous):

How many positive factors of the number 36,000,000 are not perfect squares?

OpenStudy (anonymous):

\[6^{2}\times10^{6}\]

OpenStudy (anonymous):

find the prime factorization first

OpenStudy (anonymous):

\[(2*2*3*3)*(10*10*10*10*10*10)?\]? not sure

OpenStudy (anonymous):

10 can also be broken down as 5 * 2

OpenStudy (anonymous):

yeah, so \[2^{8}*3^{3}*5^{6}\]

OpenStudy (anonymous):

In order for the number to be perfect factors, it is in the form: \[2^{2n} * 3^{2m} *5^{2p}\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

i don't fully understand... the form above?

OpenStudy (anonymous):

for 2, n can be equal to 0, 1, 2, 3, 4; 5 choices for 3, m can be equal to 0, 1; 2 choices for 5, p can be equal to 0, 1, 2, 3; 4 choices

OpenStudy (anonymous):

number of perfect squares is 5 * 2 * 4 = 40

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

9 * 4 * 7 = number of factors = 252

OpenStudy (anonymous):

number of non-perfect squares' factors = 252 - 40 = 212

OpenStudy (anonymous):

I had no clue that was how it was supposed to be done... i wish i was able to give more than a medal here, but thank you so much for your help :)

OpenStudy (anonymous):

np